### v1

@@ 6.1.3 Learning strategies
Philipp et al. [146] tackled the challenge of DG in surgical instrument localization. In this context, domain shift refers to different type of surgery, illumination, types and appearances of instruments. To overcome the problem of domain shift, the authors proposed an uncertainty-based CNN approach which adaptively chooses the most pertinent data source by integrating its own uncertainty into the inference. In particular, they trained a dynamic CNN on a single dataset which consists of combined image and optical flow modalities. The network comprises two ensemble networks, one for each modality. In order to extract the most relevant information from both data sources, the network is guided using uncertainty-based criteria which estimates the pixel-wise prediction quality for each modality. Pixel-wise uncertainty is defined as the standard deviation across the ensemble outputs. Next, the authors normalize these pixel-wise uncertainty to produce "uncertainty masks" ranging between 0 (high uncertainty) and 1 (high confidence). Finally, by weighting the predictions from each modality based on their respective uncertainties, the method ensures that the final output is influenced more by the most certain and relevant data. Then, a weighted linear combination of the predictions is computed. The proposed method was evaluated on different surgical datasets, i.e., eye and laparoscopic surgeries. It showed good generalization performances across different surgical domains.
Wang and Xia [203] proposed to use a consistency regularization loss in Domain-Ensemble Learning with Cross-Domain Mixup (DELCOM). Specifically, they consider the predicted probability vector $z_{i}^{c}$ of a domain-specific model $f_{i}$ as a pseudo label to supervise the learning of non-domain specific model $f_{j}(j\neq i)$. Given $M$ domains, an ensemble of the predictions from all non-domain-specific model is computed $z_{i}^{c(ens)}=\frac{1}{M-1}\sum_{j=1,j\neq i}^{M}z_{i}^{c(j)}$. In order to enforce the consistency between the predictions made by non-domain-specific models and the pseudo label, a consistency regularization loss is defined as follows:
$\mathcal{L}_{CR}=\sum_{i=1}^{N}||z_{i}^{c(ens)}-z_{i}^{c}||^{2}$ (13)
Each domain is selected in turn, supervised learning is applied for domain-specific model and consistency regularization learning is performed for non-domain-specific models. To enhance the consistency regularization loss, weak transformations are employed to augment the inputs of $f_{i}$ such as standard flip-and-shift transformation. In contrast, strong transformations are used to augment the inputs of ${f_{j}}_{j\neq i}$ by using a combination of RandAugment [44] and flip-and-shift transformation. This method was applied for thoracic disease classification in unseen domains and showed that it outperformed the state-of-the-art DG methods on unseen datasets.
@@
Shen et al. [171] proposed CD2-pFed, a novel Cyclic Distillation-guided Channel Decoupling framework, to personalize the global model in federated learning. To address the non-IID data across different clients, they presented a channel decoupling strategy for model personalization. The goal is to train a collection of $k$ models clients to effectively adapt to their local dataset without exchanging their data. The network of each client $i$ is composed of private personalized parameters $\theta_{i}$, and global shared parameters $\theta_{0}$. Specifically, they applied a vertical decoupling strategy consisting of assigning an adaptive proportion of learnable personalized weights at each layer from the target model, moving from the top layers to the bottom layers. A uniform personalization partition rate, ranging between zero and one, is defined to determine the precise proportion of the personalized channels in each layer. At the initial stage, it is set to a small value to facilitate learning global representation for faster convergence, and afterwards gradually increase its value. Moreover, to enhance the collaboration between private and shared weights, they used a novel Cyclic Distillation (CD) scheme to impose a consistent regularization between the local and global model representations to avoid divergence. For each input sample, they used the KL divergence to impose an consistency regularization between the $\theta_{i}$ and $\theta_{0}$ guiding the predictions from $\theta_{i}$ and $\theta_{0}$ denoted by $\hat{y_{i}}^{L}$ and $\hat{y_{i}}^{G}$ to align to each other:
$\mathcal{L}_{CD}=\frac{1}{2}(KL(\hat{y_{i}}^{L},\hat{y_{i}}^{G})+KL(\hat{y_{i}}^{L},\hat{y_{i}}^{G}))$ (14)
They tested their approach on Histo-FED, a medical image dataset, consisting of both public and private H&E stained histological WSI of human Colorectal Cancer (CRC) and normal tissue. They showed that guided by cyclical distillation, their channel decoupling framework can deliver more accurate and generalized results, outperforming the baselines.
For the purpose of training DL models robust to data heterogeneity in a federated learning setting, Andreux et al. [6] presented a SiloBN, a novel federated learning approach that uses local-statistic Batch Normalization (BN) layers. The premise of this method is that BN layers can be employed to discriminate between local and domain-invariant data. An input tensor x is normalized using a BN layer, as described in the following equation:
$BN(x)=\gamma\frac{x-\mu}{\sqrt{\sigma^{2}+\epsilon}}+\beta$ (15)
where $\gamma$ and $\beta$ are learned affine renormalization parameters along the channel axis (learned parameters), $\mu$ and $\sigma^{2}$ denote the running means and variances of each channel computed across both spatial and batch dimensions, respectively (BN statistics).
@@
In particular, the statistics from BN and its trained parameters serve distinct purposes. While the training parameters may be used across other domains, BN statistics reflect local domain peculiarities. Andreux et al. [6] proposed to take into account the separate roles of BN statistics and learned parameters. In SiloBN, only the learned BN parameters are shared across centers, whereas BN statistics remain local. This approach is considered as model personalization method as a model per center is learned, rather than one global model. To generalize the resulting models to unseen center, similar to AdaBN [116], the authors proposed to recompute BN statistics on a data batch from the target domain and keep the other model parameters frozen from the federated learning. This approach has shown promising out-of-domain generalization performances when assessed on real-world multicentric histopathology datasets.
@@
Scalbert et al. [163]  [^13] https://gitlab.com/vitadx/articles/test-time-i2i-translation-ensembling integrated TTA to their DG framework based on multi-domain image-to-image translation model StarGANv2 (which was detailed in Section 6.1.1). At test time, this method projects testing images to the source domains, classify the projected images and ensemble their predictions. This method does not rely on any prior on the domain shifts, it learns them directly from the training data. Given an unseen image, the authors sample $M$ (where $M$ is the number of source domains) random latent code vectors with respective domain label. Then, different style vectors are generated by feeding the random latent code and its domain label to a mapping network. Finally, given the testing image with the style vectors, the StarGanV2 generator translates the image to different source domains. The conducted experiments on different histopathology datasets showed that the method is more efficient than previous color augmentation/normalization, train-time data augmentation and DG methods (including CORAL[186], IRM [8], Group DRO [160] and FISH [176]).
On the other hand, recent studies have shown that test-time feature selection can achieve significant gains in performance, even when spurious correlations are learned [24]. Bissoto et al. [23] [^14] https://github.com/alceubissoto/artifact-generalization-skin proposed to use this concept in their framework. They employed test-time debiasing, where most relevant features are selected to boost the performances. More particularly, they applied NoiseCrop procedure which removes the background information, replaces it with a uniform noise, and resizes the lesion to occupy the whole image. The proposed approach was applied for skin lesion detection. For better generalization assessment, the data was partitioned into levels of increasingly higher biased training and testing sets. The results showed that their approach outperformed the baseline ERM approach and other DG methods such as RSC and GroupDRO with an AUC of 0.74 on a strong biased test set (referred to as trap test). For out-of-distribution sets, the gain of performances using debiasing depends on the similarity between the artifacts present in the testing subset and those in training, used to partition domains. Despite promising performances, test-time debiasing requires domain knowledge of the task.
@@
In the context of surgical scene understanding, Seenivasan et al. [166] proposed incremental domain generalization [^15] https://github.com/lalithjets/Domain-Generalization-for-Surgical-SceneGraph on scene graphs to predict instrument-tissue interaction during robot-assisted surgery. First, a Feature Extraction Network (FEN) ResNet18 is trained to classify the instruments and tissues in the source dataset (a nephrectomy surgery dataset). For the task of human-object interaction detection, the authors employed a graph network. After extracting the visual features of tissues and instruments using the FEN, these features are embedded in a visual graph. In parallel, nodes and node names are embedded in a semantic graph. After node aggregation in both graphs, these graphs are merged to form a combined graph $G_{c}$. Once $G_{c}$ is aggregated, its edge features are processed by a readout function to estimate interaction logits.
In order to extend the model to the target domain (transoral robotic surgery), the authors employed incremental DG which incorporates the incremental learning and the knowledge distillation. Whereas the incremental learning is used to include novel instruments in the target domain, the knowledge distillation technique is adopted to generalize the model to domain shift in the target domain. To this end, the FEN is enhanced with incremental learning to include new instruments. Since each scene is considered as a sparse graph and the graph network is robust in handling various numbers of nodes without a fixed sequence, novel instruments can be incorporated by training the graph network on the target domain data without changing the network architecture.
To handle domain shifts, the authors integrated an enhanced graph network with a teacher-student model-based knowledge distillation. In particular, the teacher network is a network trained on the source domain using the multi-label soft margin loss between the network output and the ground truth. In contrast, the student network is a copy of the teacher network. It is trained on the whole target domain dataset and on a sample of the source domain dataset using the multi-label soft margin loss. In addition, the student network is regularized using a knowledge distillation loss between the teacher and student network logits. This loss allows the student network to retain on the source domain while generalizing to the target domain. Despite promising performances, this method showed limited performances on the target domain. Thus, further improvement are required for real-world applications.
In line with the previously mentioned paper, the authors [167] designed a multi-task learning model [^16] https://github.com/lalithjets/Domain-adaptation-in-MTL to perform tool-tissue interaction detection and scene caption. The model consists of: 1) a shared feature extractor 2) a mesh-transformer branch for scene captioning and 3) a graph attention branch for tool-tissue interaction detection. To deal with domain shift, the authors proposed a class incremental contrastive learning approach for surgical scene understanding. In addition, curriculum learning is applied by using Laplacian of Gaussian (LOG) based curriculum by smoothing across all three modules to enhance model learning. In particular, curriculum by smoothing is adopted. The authors used LOG kernels, instead of the standard Gaussian kernel, to achieve better feature representation learning and highlight the instrument contours.
@@
Self-supervised learning Self-supervised learning is a paradigm in machine learning aiming to construct robust image representations via pretext tasks that do not require semantic annotations. Within the context of DG, self-supervised learning based on contrastive learning has recently emerged. In self-supervised learning, the focus shifts from relying on class labels to leveraging the structure within the data itself. Typically, it uses a pre-train strategy wherein a DL model is encouraged to develop generalized representations. It achieves this by optimizing for similarity across different augmented versions of the same image within the training dataset. These methods mostly use contrastive learning to produce pre-text invariant representations, without the supervision of annotations.
For instance, Li et al. [118] proposed to learn invariant features to various vendor-device styles by training a network with a multi-style and multi-view unsupervised self-learning scheme using contrastive learning. In their self-supervised framework [^17] https://github.com/lizheren/MSVCLMICCAI2021., an encoder is trained to learn representations using contrastive learning, without the supervision of annotations. The end goal is to pre-train a model for various downstream tasks.
In order to form the pairs for self-supervised learning, the input mini-batch ($B$) is augmented into $2B$ samples using random augmentations such as cropping, and rotation. Using the augmented mini-batch, two samples from the same image source are considered as a positive pair $(i,j)$ while the other samples within the mini-batch are regarded as negative pairs. The contrastive loss in self-supervised learning scheme, is defined as:
$\mathcal{L}_{i,j}=-log\frac{exp(sim(z_{i}^{f},z_{j}^{f})/\tau)}{\sum_{k=1}^{2B}{\mathbb{1}}_{k\neq i}exp(sim(z_{i}^{f},z_{k}^{f})/\tau)}$ (16)
where $z^{f}$ denotes the extracted features, $sim(\cdot)$ is the dot product, $\mathbb{1}_{k\neq i}\in\{0,1\}$ is an indicator function equals to 1 when $k\neq i$ and $\tau$ is a temperature parameter.
@@
Through maximizing the agreement for the positive pairs, the learned representations should be more general and robust to the augmentations or operations performed. The authors used this concept to explore the generalization for various vendor styles and the domains of CC and MLO views in mammography. More specifically, to achieve the aim of generalization to multiple vendor styles, the authors used a CycleGAN to create multiple vendor style images from a single source image. Positive pairs for multi-style contrastive learning are formed by randomly selecting two images derived from the same source image (including the original and its transformed versions). The other combinations are treated as negative pairs. In addition, they employed multi-view contrastive learning where the CC and MLO views of the same breast are considered positive views while the other combinations are negative pairs. After pre-training with contrastive loss, the backbone is used for the main task of lesion detection. The proposed method was assessed with mammograms from four vendors and one unseen public dataset. It has shown that it can help to significantly improve lesion detection on both seen and unseen domains.
Similarly, Vuong et al. [202] proposed to use a self-supervised approach using contrastive learning for achieving a domain-shift robust representation. For achieving good image representation, contrastive learning method requires to process a large batch size, which is computationally demanding, a high GPU memory is needed. To overcome this issue, Vuong et al. [202] employed Momentum Contrast (MoCo) [74] which enables the construction of a consistent dictionary of negative samples in near-linear scaling. The authors designed two dedicated momentum branch for using MoCo for both InfoMin and PatchShuffling augmentations (described in Section 6.1.1). Each branch encodes and stores a dictionary of image representations of the corresponding studied augmentation. They employed an extended version of InfoNCE loss [141], where the mutual information across positive pairs resulting from an encoder and its momentum version is maximized. In contrast, the similarity between the representation of the current view of the image and the representations of other negative samples, as encoded by the momentum encoder, is minimized.
@@
For the application of surgical scene understanding, Seenivasan et al. [167] proposed to use Class Incremental Contrastive Learning (CICL) to address intensity shift and novel class appearance in the target domain. Inspired by Xu et al. [214], they integrated supervised contrastive loss, also known as SupCon loss [98], in their framework. This technique also applies extensive augmentation to the input and maximizes the mutual information for different views. However, it also minimizes the distance between the same label inputs across domains and pushes apart the samples with different labels in the feature embedding space.
[…]
The ESL task learning aims to learn generic feature representation from a more difficult task. In ESL, images are relabeled by assigning a separate single label to images that belongs to multiple diseases. A DL model is then trained on the relabeled images using the cross-entropy loss and a class-balanced loss technique.
The $S^{4}L$ component consists of retraining the model using consistency regularization on unlabeled data. In order to enforce the model to predict with confidence on the perturbed data, they used a consistency regularization loss ($\mathcal{L}_{cr}$) for unlabeled data. It consisted of a combination of Virtual Adversarial Training (VAT) [134] loss and Conditional Entropy Minimization (CEM) [66] loss. In addition, they integrated a self-supervised learning loss where they adopted rotation pretext task. To this end, they augmented images in the dataset by creating four rotated copies from $x$ by $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$ degrees, and trained an auxiliary head to predict the rotation task.
@@
Meta-learning Meta-learning, also known as learning to learn, is a paradigm aiming to learn from episodes derived from related tasks to enhance the efficacy of future learning. The motivation for incorporating meta-learning in DG stems from the objective to expose the DL model to domain shift during training, thereby improving the model adaptability to domain shift when encountering unseen data [235].
In this context, a popular approach applying meta-learning in DG is Model-Agnostic Meta-Learning (MAML) [57]. The main idea of MAML is to transfer the knowledge from previous training tasks to facilitate the learning of new testing tasks. It uses the principle of episodic training to optimize the model according to its performance on the virtual test domain, which reflects its generalization ability. In episodic training, a simulation of domain shift is done by dividing the source domains into "meta-train" and "meta-test" domains, also called virtual training and virtual testing domains. Specifically, to enhance the generalization performance, this framework optimizes the model according to its performance on a virtual test domain. The meta-task $T_{i}$ is produced using Task Sampling (TS) strategy, i.e., it is episodically sampled from the set $T=\{T_{1},T_{2},...,T_{K}\}$. Depending on $T_{i}$, $M$ source domains are split into $M-1$ $D_{i}^{tr}$ meta-train and the held-out meta-test $D_{i}^{te}$. The model is trained using two loops of optimization. First, it is updated in inner optimization, using one or more steps of gradient descent on $D_{i}^{tr}$ with a task loss (i.e., cross-entropy loss). Second, an outer optimization is applied to evaluate the generalization performance. To be specific, in addition to the task loss, a meta-learning objective is added to emphasize certain properties among $D_{i}^{tr}$ and $D_{i}^{te}$.
Despite promising performances of MAML, Li et al. [111] argued that in clinical applications, only a few source domains are available, which may limit the capacity of training task generation and increase the risk of overfitting. To address this issue, they proposed another DG framework based on meta-learning. Their approach uses the concept of episodic training with task augmentation on medical imaging classification. To reduce task-level overfitting, the authors proposed Mixed Task Sampling (MTS) strategy which enhances the variety for the training task simulation. In comparison to the TS strategy, the meta-test domains are acquired by the interpolation among all the source domains.
Furthermore, to regularize the deep embedding of training domains, they proposed a novel "meta-objective", which regularizes the cross-domain alignment from both sample-wise and prototype-wise.
In this context, prototype refers to the centroid of the embedding for same-class samples. In particular, sample-wise alignment is used to make the features of the same class closer to each other while the features of different class far apart. The loss of sample-wise alignment uses contrastive learning and cosine distance based decision boundary. It is defined as follows:
$\begin{split}\mathcal{L}=\sum_{x_{j}^{l}\in(D_{i}^{tr},D_{i}^{te})}\{(1-Cosine(F(E(x_{j}^{l})),\varrho_{l}^{g}))\\ +\sum_{t=1}^{c}\mathbb{1}(t\neq l)Cosine(F(E(x_{j}^{l})),\varrho_{t}^{g})\}\end{split}$ (17)
where $\mathbb{1}(\cdot)$ is an indicator function, $E$ is the encoder and $F$ is a projection layer for feature extraction. $\varrho^{g}=[\varrho_{1}^{g},\varrho_{2}^{g},..,\varrho_{c}^{g}]$ is the set of domain-general prototypes, defined as the weight vectors of the classifier $C$.
In addition, they introduced domain-specific prototypes. For each class in each domain, a prototype is calculated as the centroid of the feature embeddings of samples belonging to that class. For instance, for domain $i$, the centroid of class $l$ is computed as $\varrho_{l}^{i}=\frac{1}{n_{l}^{i}}\sum_{(x_{j}^{i},y_{j}^{i}=l)\in D_{i}}F(E(x_{j}^{i}))$; where $n_{l}^{i}$ denotes the number of samples belonging to class $l$ in domain $i$. The prototypes of all classes are represented by $\varrho^{i}=[\varrho_{1}^{i},\varrho_{2}^{i},...\varrho_{c}^{i}]$. Given these prototypes, the prediction scores of an input sample are obtained by computing the cosine distance between feature embedding and prototypes. For a query sample from meta-test, $M$ prediction scores are computed based on the $M-1$ domain specific prototypes and the domain-general prototypes. Finally, a prototype-wise alignment based on KL divergence is proposed to enforce the prediction scores across different prototypes to be consistent with each other.
The proposed approach was validated on two tasks 1) epithelium-stroma classification using histopathological images and 2) liver segmentation on abdominal CT image. For the first task, it showed promising performances, outperforming the ERM baseline approach raising the accuracy from 88.43% to 91.77% on average.
@@
Bayasi et al. [18] also proposed an episodic training based on meta-learning to tackle the problem of DL model generalization. In particular, they addressed the problem of shortcut learning and feature suppression, wherein a DL model tends to only learn a subset of features and ignore the other features resulting in shortcut decision rules. They proposed BoosterNet, an auxiliary network that can be added to any arbitrary core network to enhance its generalizability without the need to change its training procedure or its architecture. Their approach combats shortcut learning using the concept of feature culpabiblity. It uses episodic learning to learn from the most culpable features (referred to as confusion features) in the core network, that is features which are linked with erroneous predictions, and to learn from the most predictive characteristics of the data, named least culpable features or discriminant features. BoosterNet was validated for detecting skin lesions, where the authors showed improved generalization performance compared to other benchmark DG approaches including data augmentation based DG, adversarial training, and feature alignment.
Sikaroudi et al. [178] studied a Model Agnostic Learning of Semantic Features (MASF) strategy in DG. MASF was inspired by the MAML approach, it aims to learn a latent space representation suitable for generalization to an unseen test domain. The architecture consists of layers of feature extractor, classification and metric embedding.
MASF also uses episodic training process. In each iteration, a batch of the source domain is split into meta-train and meta-test batches. The DL model is trained with the cross-entropy loss for supervised learning and hard class separation. In addition, for enhancing soft class separation and promoting domain-independent class-specific cohesion and separation of instance, a metric loss is added. Specifically, the metric loss is the average triplet loss for a batch of triplets, which is formed from an anchor, positive and negative instances from all the source domain dataset. The triplet loss compares a reference input (referred to as anchor, $x_{A}^{b}$) to a matching input (positive instance belonging to the same class as $x_{A}^{b}$, denoted by $x_{P}^{b}$) and a non-matching input (negative instance belonging to a different class than $x_{A}^{b}$, denoted by $x_{N}^{b}$). The distance from the anchor to the positive instances is minimized, while the distance form the anchor to the negative instances is maximized. The triplet loss is formulated as follows:
$\begin{split}\mathcal{L}_{triplet}&=\frac{1}{B}\sum_{b=1}^{B}max(||F(E(x_{A}^{b}))-F(E(x_{P}^{b}))||_{2}^{2}-\\ &||F(E(x_{A}^{b}))-F(E(x_{P}^{b}))||_{2}^{2}+\alpha,0)\end{split}$ (18)
where $||\cdot||_{2}$ is the $l_{2}$ norm, $\alpha$ is a margin between positive and negative pairs, $B$ denotes number of instances in a batch, $E$ denotes the sub-network feature extractor that outputs a latent space representation, and $F$ is a projection layer which lower the dimension of the representation layer.
Finally, to enforce representation alignment, the authors employed the KL divergence. Given two source domains $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, this loss can be formulated as follows:
$\mathcal{L}_{alignment}(\mathcal{S}_{i},\mathcal{S}_{j})=\frac{1}{c}\sum_{k=1}^{c}\frac{1}{2}[D_{KL}(s_{k}^{(i)}||s_{k}^{(j)})+D_{KL}(s_{k}^{(j)}||s_{c}^{(i)})]$ (19)
where $D_{KL}$ represents the symmetrized Kullback-Leibler divergence, $s_{k}^{(.)}$ denotes the soft confusion matrix, which is computed by applying softmax function with temperature $\tau>1$ on the output of classification network.
The training process consists of first updating the weights of the classifier and the feature extractor using the cross-entropy loss. Then, the meta loss is computed using a weighted sum of the alignment loss ($\mathcal{L}_{alignment}$) and the triplet loss ($\mathcal{L}_{triplet}$). It is combined to the cross-entropy loss. Next, two gradient descent steps are done to update the network according to 1) the meta loss and the cross-entropy loss and 2) to the triplet loss.
They applied this method for Renal Cell Carcinoma (RCC) subtypes classification in WSI. The results showed that this method outperformed the baseline, which was training on only cross-entropy loss on three hold-out trial sites.

### v3

@@ 5.1.3 Learning strategies
In the context of surgical instrument localization, Philipp et al. [56] proposed an uncertainty-based dynamic CNN which combines two modalities (image and optic flow modality). Their CNN dynamics were guided using pixel level uncertainty estimated separately for each modality. It comprises two ensemble network, one for each modality. The outputs for each ensemble is fused using the mean of the prediction maps. Next, pixel-wise uncertainty map is estimated using the standard deviation across the ensemble individuals. Uncertainty masks are then computed by normalizing these pixel-wise uncertainty maps. Finally, these uncertainty masks are used to fuse the ensemble for the two modalities by weighting the predictions from each modality based on their respective uncertainties. This method showed good generalization performances when evaluated on heterogeneous surgical datasets coming from different domains including eye, laparoscopic and neurosurgeries.
Wang and Xia [35] proposed domain-ensemble learning with cross-domain mixup. Their model comprised a shared backbone for all source domains and a domain-specific classifier. After training their domain specific model, they used ensemble learning to expose the model optimization to domain distribution discrepancy. They enforced the consistency between the predictions obtained by all non-domain specific model (ensemble of predictions) and a pseudo label generated by a domain-specific model (i.e., prediction from a domain-specific model). This method was applied for thoracic disease classification in unseen domains and showed that it outperformed the state-of-the-art DG methods on unseen datasets.
@@
In their Federated Learning (FL) framework, Shen et al. [57] and Andreux et al. [58] addressed the non-IID data across different clients. Shen et al. [57] presented a channel decoupling strategy for model personalization. The network of each client ($i$) was composed of private personalized parameters $\theta_{i}$, and global shared parameters $\theta_{0}$. Their vertical decoupling strategy consisted of assigning an adaptive proportion of learnable personalized weights at each layer from the target model, moving from the top layers to the bottom layers. A uniform personalization partition rate, ranging between zero and one, was defined to determine the precise proportion of the personalized channels in each layer. To enhance the collaboration between private and shared weights, they used a cyclic distillation scheme. For each input sample, they used the KL divergence to impose a consistency regularization between $\theta_{i}$ and $\theta_{0}$, guiding the predictions from $\theta_{i}$ and $\theta_{0}$ to align to each other. They showed that their channel decoupling framework can deliver more accurate and generalized results, outperforming the baselines when evaluated on Histo-FED dataset.
@@
Andreux et al. [58] presented SiloBN, another model personalization method based on FL. It uses local-statistic Batch Normalization (BN) layers to discriminate between local and domain-invariant data. Only the learned BN parameters are shared across centers, whereas BN statistics (the running means and variances of each channel computed across both spatial and batch dimensions, respectively) remain local. To generalize the resulting models to unseen centers, similar to AdaBN [133], the BN statistics are recomputed on a data batch from the target domain while the other model parameters are kept frozen as obtained from the federated training. This approach has shown promising out-of-domain generalization performances when assessed on real-world multicentric histopathology datasets.
@@
Scalbert et al. [41] [^13] https://gitlab.com/vitadx/articles/test-time-i2i-translation-ensembling integrated TTA to their DG framework based on StarGANv2. At test time, this method projects testing images to $M$ (where $M$ is the number of source domains) source domains, classify the projected images and ensemble their predictions. Given an unseen image, $M$ style vectors are first encoded by feeding a random latent code and its domain label to a mapping network. The StarGANV2 generator then takes the style vector and the testing image as input to translate the image to different source domains. Experiments on different histopathology datasets showed that this method is more efficient than previous color augmentation/normalization, train-time data augmentation and DG methods.
Bissoto et al. [59] [^14] https://github.com/alceubissoto/artifact-generalization-skin proposed test-time debiasing where feature selection is performed during inference. The idea is to force the network to use the correct correlations learned to make the prediction. To reduce spurious features in testing images, NoiseCrop was applied. It removes the background information, replaces it with a uniform noise, and resizes the lesion to occupy the whole image. For the task of skin lesion detection, this approach outperformed the baseline ERM approach and other DG methods such as RSC and GroupDRO with an AUC of 0.74 on a strong biased test set. Despite promising performances, test-time debiasing requires domain knowledge of the task.
@@
Seenivasan et al. [60] proposed incremental DG [^15] https://github.com/lalithjets/Domain-Generalization-for-Surgical-Scene-Graph on scene graphs to predict instrument-tissue interaction during robot-assisted surgery. They trained a feature extraction network and a graph network on a nephrectomy surgery dataset to classify 9 classes. The feature extraction network was then extended to the target domain (a transoral robotic surgery dataset) to classify 11 classes using an incremental learning technique, as described in [134]. In addition, the authors proposed to use knowledge distillation, where the teacher network is a network trained on the source domain and the student network is a copy of the teacher network, trained on the whole target domain dataset and on a sample of the source domain dataset. To further enable the student network to retain the knowledge from the source domain while generalizing to the target domain, it was regularized using a knowledge distillation loss between the teacher and student network logits. Despite promising performances, this method showed limited performances on the target domain.
In line with the previously mentioned paper, the authors [61] designed a multi-task learning model [^16] https://github.com/lalithjets/Domain-adaptation-in-MTL to perform tool-tissue interaction detection and scene caption. The model consists of: 1) a shared feature extractor 2) a mesh-transformer branch for scene captioning and 3) a graph attention branch for tool-tissue interaction detection. To deal with domain shift, the authors proposed a class incremental contrastive learning approach for surgical scene understanding. In addition, they developed Laplacian of Gaussian (LOG) based curriculum by smoothing across all three modules to enhance model learning. This approach used LOG kernels instead of Gaussian kernel to control the features entering the model at the initial epochs and highlight the instrument contours, thus allowing the model to learn gradually.
@@
Self-supervised learning Self-supervised Learning (SSL) aims to construct robust image representations via pretext tasks that do not require semantic annotations, leveraging the structure within the data itself. Within the context of DG, SSL based on Contrastive Learning (CL) has recently emerged as a pre-training strategy to produce generalized, pre-text invariant representations. The pre-trained model can then be adopted for various downstream tasks.
@@
For instance, to achieve invariant representations in their SSL framework [^17] https://github.com/lizheren/MSVCL_MICCAI2021, Li et al. [39] employed two types of CL: 1) a multi-style CL to generalize to multiple device style and 2) a multi-view CL to learn representations that are robust to the CC and MLO views in mammography. For multi-style CL, a CycleGAN was utilized to create multiple device-style images from a single source image. Positive pairs were formed by randomly selecting two images derived from the same source image. For multi-view CL, the CC and MLO views of the same breast were considered positive pairs. Following the pre-training stage, the backbone was used for the main task of lesion detection. The proposed method was assessed with mammograms from four vendors and one unseen public dataset (INbreast). It has shown significant improvement for lesion detection on both seen and unseen domains.
@@
For the application of surgical scene understanding, Seenivasan et al. [61] proposed to use a hybrid approach combining self-supervised learning scheme and supervised learning. Inspired by Xu et al. [135], they integrated supervised contrastive loss, also known as SupCon loss [136]. Similar to self-supervised contrastive learning, this technique applies extensive augmentation to the input and maximizes the mutual information for different views. However, it also leverages the label information: it minimizes the distance between the same label inputs across domains and pushes apart the samples with different labels in the feature embedding space.
Vuong et al. [43] employed Momentum Contrast (MoCo) [137], which uses CL as dictionary look-up: an encoded “query” (image) should be similar to its matching key and dissimilar to others. In MoCo, a dynamic dictionary is implemented with a queue and a moving-averaged encoder (momentum encoder). The dictionary keys are defined on-the-fly by a set of data samples and are encoded by a momentum encoder. Vuong et al. [43] used this concept and designed two dedicated momentum branches for both InfoMin and PatchShuffling augmentations. Each branch encodes and stores a dictionary of image representations for the corresponding augmentation. The network was optimized using an extended version of InfoNCE loss [138]. When evaluated on unseen dataset for colorectal cancer tissue classification, this approach outperformed other SSL methods such as MoCoV2 which uses a single momentum branch.
@@
Meta-learning Meta-learning, also known as learning to learn, is a paradigm aiming to learn from episodes derived from related tasks to enhance the efficacy of future learning. It has been applied to DG, by adopting an episodic training paradigm, where at each iteration a meta-task is generated with the source domains splitted into meta-train and meta-test domains to simulate domain shift.
To address the problem of DG with limited data, Li et al. [63] proposed a mixed task sampling strategy where the meta-test domains were generated by interpolating among all the source domains. In their meta-objective, a regularization was incorporated to enforce the alignment of embeddings across training domains from both sample-wise and prototype-wise perspectives. Sample-wise alignment reduces intra-class distances while increasing inter-class separations using CL and cosine distance based loss. Domain-general prototypes were the weight vectors of the classifier, and domain-specific prototypes were the centroid of the embedding for same-class samples for each domain. A prototype-wise alignment based on KL divergence was proposed to enforce the prediction scores across different prototypes to be consistent with each other. This approach outperformed the ERM baseline approach raising the accuracy from 88.43% to 91.77% on average for epithelium-stroma classification using histopathological images.
@@
Bayasi et al. [64] proposed BoosterNet, an auxiliary network that can be added to any arbitrary core network to enhance its generalizability without the need to change its training procedure or its architecture. Their approach combats shortcut learning using the concept of feature culpabiblity. It uses episodic learning to learn from the most culpable features in the core network (i.e., features which are linked with erroneous predictions) and from the most predictive characteristics of the data (discriminant features). BoosterNet was validated for detecting skin lesions, where it showed improved generalization performance compared to other benchmark DG approaches including data augmentation based DG, adversarial training, and feature alignment.
Inspired by Model Agnostic Learning of Semantic Features (MASF), Sikaroudi et al. [65] proposed to learn a latent space representation suitable for generalization to an unseen test domain. Their meta-objective was a weighted sum of an alignment loss and a triplet loss. The alignment loss was the KL divergence between the soft confusion matrix of different domains. The metric loss was the average triplet loss for a batch of triplets, which is formed from an anchor, positive and negative instances from all the source domain dataset. The triplet loss compares a reference input (an anchor) to a matching input (positive instance belonging to the same class as the anchor) and a non-matching input (negative instance belonging to a different class than the anchor). The distance from the anchor to the positive instances is minimized, while the distance form the anchor to the negative instances is maximized. For the task of renal cell carcinoma subtypes classification in WSI, this method outperformed the baseline, which involved training using only cross-entropy loss on three hold-out trial sites.